OpenStudy (anonymous):

–5y – 9 = –(y – 1) I don't know how but I got 3/10 but thats not one of the multiple choice answers lol

OpenStudy (anonymous):

Your answer is fall Modom :( Do u want my help ?

OpenStudy (uri):

-5y-9=-y+1 Add 9 both sides -5y-9=-y+1 9 9 ========= -5y=-y+!0 Add y both sides -5y=-y+!0 y y ========= -4y=10 DIvide 10 with -4 y=10/4 y=5/2

OpenStudy (anonymous):

@vt : Got it ?????

OpenStudy (anonymous):

Ohhhh Lmao I was way off, thank you : )

OpenStudy (anonymous):

heh :) Your welcome Modom:)

OpenStudy (uri):

Gimmee medalll @vt

OpenStudy (anonymous):

Jajaja ok I gotchaa